1.)How much energy is lost by 20 kg copper if it is cooled from 50℃ to 30℃?



2.)A 20 kg metal rod is heated 100oC with 2000 J of energy. What is the specific heat of the metal?

1
Expert's answer
2022-01-10T13:48:26-0500

1)

"Q=mc\Delta T,""Q=20\ kg\times385\ \dfrac{J}{kg\times\!^{\circ}C}\times(50^{\circ}C-30^{\circ}C),""Q=154000\ J=154\ kJ."


2)

"Q=mc\Delta T,""c=\dfrac{Q}{m\Delta T},""c=\dfrac{2000\ J}{20\ kg\times100^{\circ}C}=1.0\ \dfrac{J}{kg\times\!^{\circ}C}."